OpenStudy (anonymous):

Find the area of a triangle of a rectangle with a length of 6 cm and width of 7 cm. Do not include units in your answer.

OpenStudy (anonymous):

L*w=A so 6*7=A=42=(b*h/2) b=7 solve for h=12 :)

OpenStudy (anonymous):

in order to solve for do you divide 12 by 6?

OpenStudy (anonymous):

um nope :P

OpenStudy (anonymous):

6 is irrilvant for the triangle :)

OpenStudy (anonymous):

so what do you use to solve for 12 than?

OpenStudy (asnaseer):

is this what you mean?  i.e. find the area of one of these triangles? if yes, then area equals half the area of the rectangle itself.

OpenStudy (anonymous):

it doesnt show a picture but i think the answer is 2